# xbx-supp-arxiv-0

### v1

@@ 3.3 Maximum approximate likelihood
From equation (9), the log-likelihood function of the XBX regression model can be approximated using the Gauss-Laguerre quadrature rule. The parameters are the regression coefficients $\beta$ and $\gamma$ from (2) as well as the log exceedance parameter $\xi=\log\nu$, to assure that all parameters have unrestricted ranges. This yields the following approximate log-likelihood
@@
where $\psi(k)$ is the digamma function, and ${}_{3}F_{2}$ is the generalized hypergeometric function.
@@ 3.4 Inference
$2\left\{\ell^{(a)}(\hat{\beta},\hat{\gamma},\hat{\xi})-\max_{C\theta=b}\ell^{(a)}(\beta,\gamma,\xi)\right\}\,,$
@@
$(R\hat{\theta}-b)^{\top}\{RI^{(a)}(\hat{\theta})^{-1}R^{\top}\}(R\hat{\theta}-b)\,,$
@@ 4.2 Data
Interest is in linking the proportion $Y_{i}\in[0,1]$ of the total tokens invested by the $i$th subject in the nine rounds out of the $900$ available tokens to explanatory information. The explanatory information we consider here consists of the subject characteristics and their interactions that Glätzle-Rützler et al. (2015) found to be relevant in the experiment. These are Grade ($G_{i}$ with value 0 and 1, according to whether the the $i$th subject is from lower grades 6–8 or upper grades 10–12, respectively), Arrangement ($T_{i}$, with values 0 and 1, according to whether the $i$th subject is an individual or a team, respectively), Age ($A_{i}$ the age of the player or the average age of the team of players in years), Sex ($S_{i}$, with value 1 if the player was male, or the team included at least one male student, and 0, otherwise), and the interactions of Grade with Arrangement and Age.
@@ 4.3 Models
The first model is a normal linear model (N) under which $Y_{i}$ has a normal distribution with mean $\mu_{i}=\eta_{i}$ and variance $\sigma_{i}=\sigma^{2}$. This model corresponds to the OLS approach used by Glätzle-Rützler et al. (2015). The second model we consider is the heteroscedastic two-limit tobit model, which is based on the censored normal distribution (CN) and described in Section 1.2, and where the latent variable $Y_{i}^{*}$ has mean $\mu_{i}=\eta_{i}$ and variance $\sigma_{i}$ satisfying $\log(\sigma_{i})=\zeta_{i}$ with
@@ 4.4 Estimates
Coefficient estimates (and standard errors) along with log-likelihood, Akaike and Bayes information criteria (AIC and BIC) for all four models are provided in Table 1 and Table 2, respectively. Separate tables are used because the coefficients from N and CN use an identity link for the mean parameter whereas B and XBX use a logit link. In addition, the log-likelihood, and, hence, AIC and BIC, are comparable only between CN and XBX because those two models have the same support for the response variable, that is the unit interval with point masses at 0 and 1.
@@
Figure 4: Hanging rootograms for the fitted models in Table 1 and Table 2. The rootograms compare the histogram estimate of the marginal distribution of the proportions of tokens invested with the corresponding expected frequencies from the fitted models. The square root of the expected frequencies are shown as red dots and the square root of the observed frequencies are hanging from the points as gray bars. The quality of the fit is determined by checking whether the bars hang down to the horizontal line at zero or whether they are clearly too large or too small. The dashed lines are the Tukey warning limits at ±$\pm 1$ (see Kleiber and Zeileis, 2016, for details).
@@ 4.5 Assessment of fitted models
The differences between the models become more apparent when considering the fitted distributions. Figure 4 shows hanging rootograms (see Kleiber and Zeileis, 2016, for details) for the fitted models in Table 1 and Table 2. The rootograms compare the empirical marginal distributions of the proportion of tokens invested to the implied distributions from the fitted models. The quality of the fit is determined by checking whether the bars hang down to the horizontal line at zero or whether they are clearly too large or too small. As is apparent, models N and B fit poorly in the tails. For N the support of the fitted distributions is the real line and, as a result, the expected frequencies in intervals close to 00 or $1$ are larger and smaller, respectively, than what is observed. The corresponding expected frequencies for B are smaller than what is observed, because beta regression struggles with giving enough probability mass to intervals close to 00 or $1$ because it cannot assign point masses at the boundaries. Also, the expected frequencies at intervals around 0.5 are too small compared to what is observed. In contrast, models CN and XBX fit very well with almost all bars hanging close to the zero reference line. The fit for XBX appears to be slightly better which is also reflected by both AIC and BIC being smaller for XBX; see Table 1 and Table 2.
@@ 4.7 Three-part hurdle models
In our analysis, we have not considered the three-part hurdle model of Ospina and Ferrari (2010) (“zero-and-one-inflated” beta) for two reasons: (a) from an economic point of view, it is not plausible that the effects driving “fully rational” behavior ($Y=1$) are different from the effects driving “very rational” behavior (say, $Y<0.95$ or $Y=1$); (b) the hurdle needs substantially more parameters than XBX due to the regression structures for modeling the probability of boundary and non-boundary observations, whose estimation may also not be reliable. For example, for the loss aversion data, employing a multinomial logistic regression model for the probabilities of boundary and non-boundary observations with the same covariates as those in the linear predictor (16) or (17), and model B for the non-boundary observations, results in a three-part hurdle model with 25 or 19 parameters, respectively, instead of the 12 parameters that XBX requires. Notably, maximum likelihood estimation results in infinite estimates for the parameters of either multinomial logistic regression model (see Table LABEL:tab:lossaversion-beta01-1 and Table LABEL:tab:lossaversion-beta01-2 in the supplementary material document for estimates and estimated standard errors) due to data separation. Also, the BIC values from both three-part hurdle model are larger than those of the CN and XBX models.
@@ 5 XBX versus heteroscedastic two-limit tobit regression
Theorem 1 shows that regression models based on the XB distribution allow to parameterize the transition from the beta distribution to the censored normal distribution with a single parameter $u$. On the other hand, XBX regression is a continuous mixture of XB distributions that shrinks towards a beta regression model. Hence, a comparison of XBX regression with heteroscedastic two-limit tobit regression can be based on data sets from the XB distributions. For a given $(\beta_{1},\beta_{2},\gamma_{1},\gamma_{2},u)^{\top}$, we let $\log\{\mu_{i}/(1-\mu_{i})\}=\beta_{1}+\beta_{2}x_{i}$ and $\log\phi_{i}=\gamma_{1}+\gamma_{2}x_{i}$, where $x_{i}=2(i-1)/(n-1)-1\in[-1,1]$ $(i=1,\ldots,n)$, and simulate response values $y_{1},\ldots,y_{n},y_{1}^{\prime},\ldots,y_{n}^{\prime}$ as realizations of independent random variables $Y_{1},\ldots,Y_{n},Y_{1}^{\prime},\ldots,Y_{n}^{\prime}$ with $Y_{i}\sim{\rm XB}(\mu_{i},\phi_{i},u)$, and $Y_{i}^{\prime}\sim{\rm XB}(\mu_{i},\phi_{i},u)$. We then fit the XBX and CN models on $y_{1},\ldots,y_{n}$ with covariates $x_{1},\ldots,x_{n}$. The predictive performance of each model is evaluated as $S=\sum_{i=1}^{n}C(\hat{F}_{i},y_{i}^{\prime})$, where $\hat{F}_{i}(\cdot)$ is the model’s cumulative distribution function at $x_{i}$ at the parameter estimates from $y_{1},\ldots,y_{n}$, and
@@
is the continuous ranked probability score (CRPS; Gneiting and Raftery, 2007), which is a strictly proper scoring rule.
@@
Because $x_{1},\ldots,x_{n}$ define an equispaced grid of size $n$ on $[-1,1]$, the parameters $\beta_{0},\beta_{1},\gamma_{0},\gamma_{1}$ are uniquely determined by $\mu_{1}$, $\mu_{n}$, $\phi_{1}$, and $\phi_{n}$. Also, for $\mu_{1}\leq\mu_{n}$ and $\phi_{1}\leq\phi_{n}$, it necessarily holds that $\mu_{i}\in[\mu_{1},\mu_{n}]$ and $\phi_{i}\in[\phi_{1},\phi_{n}]$. Table 4 shows the intervals $[\mu_{1},\mu_{n}]$ and $[\phi_{1},\phi_{n}]$ we consider in the simulation experiment, along with the implied parameter values. The intervals cover narrow and wide ranges for $\mu$, being either close to the boundary values or around 0.5, and narrow and wide intervals for $\phi_{i}$, with both low and high values of the precision parameters.
@@
We set $n=500$. In order to avoid trivial response configurations, from the possible combinations of $u\in\{2^{-6},2^{-5},\ldots,2\}$, and $[\mu_{1},\mu_{n}]$ and $[\phi_{1},\phi_{n}]$ in Table 4, we consider those for which the average probability of non-boundary response values in $(0,1)$ is more than $0.05$. We then simulate $100$ independent vectors $(y_{1},\ldots,y_{n},y_{1}^{\prime},\ldots,y_{n}^{\prime})$ per parameter setting, and compute the relative change in $S$, $S_{\rm CN}/S_{\rm XBX}-1$, when moving from the XBX to the CN model, based on their maximum likelihood fits. Positive values of the relative change correspond to $S_{\rm CN}>S_{\rm XBX}$, and, thus, worse predictive performance of the CN fit from the heteroscedastic two-limit tobit model relative to XBX regression.
@@
Figure 6 shows the average relative change for a subset of the parameter settings we consider; see, Figure LABEL:fig-supp:xbx-vs-htobit in the supplementary material document, for all settings. The XBX model is performing the same or outperforms the CN model in the majority of settings. CN is found to start performing slightly better for moderate values of $u$ when the response distributions have small variance (large values of $\phi$). The latter settings produce datasets where the relationship between the non-boundary responses and covariates is well-modeled as linear; see Figure LABEL:fig-supp:xb-data-1 to Figure LABEL:fig-supp:xb-data-4 in the supplementary material document show simulated data sets for all parameter settings. The two models start performing similarly again for large $u$, when the response values start concentrating on the boundaries.
@@
Figure 6: Average relative change in $S$ when moving from XBX to CN, based on $100$ samples of size $n=500$ for each combination of $u\in\{2^{-6},2^{-5},\ldots,2\}$, and a subset of $[\mu_{1},\mu_{n}]$ (columns) and $[\phi_{1},\phi_{n}]$ (rows) intervals in Table 4. The dotted vertical lines are the smallest values of $u$ for which the average probability of boundary observations exceeds $0.95$ for each combination of $[\mu_{1},\mu_{n}]$ and $[\phi_{1},\phi_{n}]$.
@@ 6 Concluding remarks
We have introduced XBX regression for modeling bounded responses with or without boundary observations. At the core of the XBX regression model is the XB distribution, which is a censored version of a four-parameter beta distribution with the same exceedance $u$ on the left and right of $(0,1)$. We prove that both beta regression with dispersion effects and the heteroscedastic two-limit tobit model are special cases of XB regression, depending on whether $u$ is zero or infinity, respectively. To avoid identifiability issues that may arise in estimating $u$ due to the similarity of the beta and normal distribution for specific parameter settings that result in low skewness and kurtosis, we define XBX regression as a shrunken version of XB regression towards beta regression, by letting $u$ have an exponential distribution with mean $\nu$. The associated marginal likelihood can be conveniently and accurately approximated using a Gauss-Laguerre quadrature rule, resulting in efficient estimation and inference procedures. In Section 5, we produce evidence that XBX regression can substantially outperform heteroscedastic two-limit tobit regression when the common features of skewness and kurtosis of bounded-domain variables are encountered, and perform only slightly worse when skewness and kurtosis are small.
@@
Finally, Bayesian inference for XBX regression is straightforward through template modeling frameworks such as JAGS (Plummer, 2003) and Stan (Carpenter et al., 2017), due to the alternative definition of XBX regression in Section 3 as the censored version of a mixture of four-parameter beta regressions with the same exponentially-distributed exceedance on the left and right of $(0,1)$.

### v2

@@ 3.3 Maximum approximate likelihood
From equation (9), the log-likelihood function of the XBX regression model can be approximated using the Gauss-Laguerre quadrature rule. The parameters are the regression coefficients $\beta$ and $\gamma$ from (2) as well as the log exceedance parameter $\xi=\log\nu$, to assure that all parameters have unrestricted ranges. This yields the approximate log-likelihood
@@
where $\psi(\cdot)$ is the digamma function, and ${}_{3}F_{2}(\cdot,\cdot,\cdot\,;\,\cdot,\cdot)$ is the generalized hypergeometric function.
@@ 3.4 Inference
$2\left\{\ell^{(a)}(\hat{\beta},\hat{\gamma},\hat{\xi})-\max_{R\theta=b}\ell^{(a)}(\beta,\gamma,\xi)\right\}\,,$
@@
$(R\hat{\theta}-b)^{\top}\{RI^{(a)}(\hat{\theta})^{-1}R^{\top}\}^{-1}(R\hat{\theta}-b)\,,$
@@ 4.2 Data
Interest is in linking the proportion $Y_{i}\in[0,1]$ of the total tokens invested by the $i$th subject in the nine rounds out of the $900$ available tokens to explanatory information. The explanatory information we consider here consists of the subject characteristics and their interactions that Glätzle-Rützler et al. (2015) found to be relevant in the experiment. These are Grade ($G_{i}$ with value 0 and 1, according to whether the the $i$th subject is from lower grades 6–8 or upper grades 10–12, respectively), Arrangement ($T_{i}$, with values 0 and 1, according to whether the $i$th subject is an individual or a team of two, respectively), Age ($A_{i}$ the age of the player or the average age of the team of players in years), Sex ($S_{i}$, with value 1 if the player was male, or the team included at least one male student, and 0, otherwise), and the interactions of Grade with Arrangement and Age.
@@ 4.3 Models
The first model is a normal linear regression model (N) under which $Y_{i}$ has a normal distribution with mean $\mu_{i}=\eta_{i}$ and variance $\sigma_{i}=\sigma^{2}$. This model corresponds to the OLS approach used by Glätzle-Rützler et al. (2015). The second model we consider is the heteroscedastic two-limit tobit model, which is based on the censored normal distribution (CN) and described in Section 1.2, and where the latent variable $Y_{i}^{*}$ has mean $\mu_{i}=\eta_{i}$ and variance $\sigma_{i}$ satisfying $\log(\sigma_{i})=\zeta_{i}$ with
@@ 4.4 Estimates
Coefficient estimates (and estimated standard errors) along with log-likelihood, Akaike and Bayes information criteria (AIC and BIC) for all four models are provided in Table 1 and Table 2, respectively. Separate tables are used because the coefficients from N and CN use an identity link for the mean parameter whereas B and XBX use a logit link. In addition, the log-likelihood, and, hence, AIC and BIC, are comparable only between CN and XBX because those two models have the same support for the response variable, that is the unit interval with point masses at 0 and 1.
@@
Figure 4: Hanging rootograms for the fitted models in Table 1 and Table 2. The rootograms compare the histogram estimate of the marginal distribution of the proportions of tokens invested with the corresponding expected frequencies from the fitted models. The square root of the expected frequencies are shown as dots connected by a solid line (in black) and the square root of the observed frequencies are hanging from the points as grey bars. The quality of the fit is determined by checking whether the bars hang down to the horizontal line at zero or whether they are clearly too large or too small. The dashed horizontal lines are the Tukey warning limits at ±$\pm 1$ (see Kleiber and Zeileis, 2016, for details).
@@ 4.5 Assessment of fitted models
The differences between the models become more apparent when considering the fitted distributions. Figure 4 shows hanging rootograms (see Kleiber and Zeileis, 2016, for details) for the fitted models in Table 1 and Table 2. The rootograms compare the empirical marginal distributions of the proportion of tokens invested to the implied distributions from the fitted models. The quality of the fit is determined by checking whether the bars hang down to the horizontal line at zero or whether they are clearly too large or too small. As is apparent, models N and B fit poorly in the tails. For N the support of the fitted distributions is the real line and, as a result, the expected frequencies in intervals close to 00 or $1$ are larger and smaller, respectively, than what is observed. The corresponding expected frequencies for B are larger than what is observed, while the expected frequencies at intervals around 0.5 are too small compared to what is observed. In contrast, models CN and XBX fit very well with all but the second bar from the left hanging close to the zero reference line and within the Tukey limits. The misfit of the second bar is larger for CN compared to XBX, which is also reflected by both AIC and BIC being smaller for XBX; see Table 1 and Table 2.
@@ 4.7 Three-part hurdle models
In our analysis, we have not considered the three-part hurdle model of Ospina and Ferrari (2010) (“zero-and-one-inflated” beta) for two reasons: (a) from an economic point of view, it is not plausible that the effects driving “fully rational” behavior ($Y=1$) are different from the effects driving “very rational” behavior (say, $Y<0.95$ or $Y=1$); (b) the hurdle needs substantially more parameters than XBX due to the regression structures for modelling the probability of boundary and non-boundary observations, whose estimation may also not be reliable. For example, for the loss aversion data, employing a multinomial logistic regression model for the probabilities of boundary and non-boundary observations with the same covariates as those in the linear predictor (16) or (17), and model B for the non-boundary observations, results in a three-part hurdle model with 25 or 19 parameters, respectively, instead of the 12 parameters that XBX requires. Notably, maximum likelihood estimation results in infinite estimates for the parameters of either multinomial logistic regression model (see Table LABEL:tab:lossaversion-beta01-1 and Table LABEL:tab:lossaversion-beta01-2 in the supplementary material document for estimates and estimated standard errors) due to data separation. Also, the BIC values from both three-part hurdle model are larger than those of the CN and XBX models.
@@ 5 XBX versus heteroscedastic two-limit tobit regression
Theorem 1 shows that regression models based on the XB distribution allow to parameterize the transition from the beta distribution to the censored normal distribution with a single parameter $u$. On the other hand, XBX regression is a continuous mixture of XB distributions that shrinks towards a beta regression model. Hence, a comparison of XBX regression with heteroscedastic two-limit tobit regression can be based on data sets from the XB distributions. For a given $(\beta_{1},\beta_{2},\gamma_{1},\gamma_{2},u)^{\top}$, we let $\log\{\mu_{i}/(1-\mu_{i})\}=\beta_{1}+\beta_{2}x_{i}$ and $\log\phi_{i}=\gamma_{1}+\gamma_{2}x_{i}$, where $x_{i}=2(i-1)/(n-1)-1\in[-1,1]$ $(i=1,\ldots,n)$, and simulate response values $y_{1},\ldots,y_{n},y_{1}^{\prime},\ldots,y_{n}^{\prime}$ as realizations of independent random variables $Y_{1},\ldots,Y_{n},Y_{1}^{\prime},\ldots,Y_{n}^{\prime}$ with $Y_{i}\sim{\rm XB}(\mu_{i},\phi_{i},u)$, and $Y_{i}^{\prime}\sim{\rm XB}(\mu_{i},\phi_{i},u)$. We then fit the XBX and CN models on $y_{1},\ldots,y_{n}$ with covariates $x_{1},\ldots,x_{n}$ in both mean and variance and precision regression specifications. The predictive performance of each model is evaluated as
$S=\sum_{i=1}^{n}C(\hat{F}_{i},y_{i}^{\prime})\,,$ (18)
@@
Because $x_{1},\ldots,x_{n}$ define an equispaced grid of size $n$ on $[-1,1]$, the parameters $\beta_{1},\beta_{2},\gamma_{1},\gamma_{2}$ are uniquely determined by $\mu_{1}$, $\mu_{n}$, $\phi_{1}$, and $\phi_{n}$. Also, for $\mu_{1}\leq\mu_{n}$ and $\phi_{1}\leq\phi_{n}$, it necessarily holds that $\mu_{i}\in[\mu_{1},\mu_{n}]$ and $\phi_{i}\in[\phi_{1},\phi_{n}]$. Table 4 shows the intervals $[\mu_{1},\mu_{n}]$ and $[\phi_{1},\phi_{n}]$ we consider in the simulation experiment, along with the implied parameter values. The intervals cover narrow and wide ranges for $\mu$, being either close to the boundary values or around 0.5, and narrow and wide intervals for $\phi_{i}$, with both low and high values of the precision parameters.
@@
We set $n=500$. In order to avoid trivial response configurations, from the possible combinations of $u\in\{2^{-6},2^{-5},\ldots,2\}$, and $[\mu_{1},\mu_{n}]$ and $[\phi_{1},\phi_{n}]$ in Table 4, we consider those for which the average probability of non-boundary response values in $(0,1)$ is more than $0.05$. We then simulate $100$ independent vectors $(y_{1},\ldots,y_{n},y_{1}^{\prime},\ldots,y_{n}^{\prime})$ per parameter setting, and compute the relative change in $S$, $S_{\rm CN}/S_{\rm XBX}-1$, when moving from the XBX to the CN model, based on their maximum likelihood fits. Positive values of the relative change correspond to $S_{\rm CN}>S_{\rm XBX}$, and, thus, worse predictive performance of the CN model relative to the XBX regression model.
@@
Figure 6 shows the average relative change for a subset of the parameter settings we consider; see, Figure LABEL:fig-supp:xbx-vs-htobit in the supplementary material document, for all settings. The XBX model is performing the same or outperforms the CN model in the majority of settings. CN is found to perform slightly better when the response distributions have small variance (large values of $\phi$) for moderate values of $u$. Those settings produce datasets where the relationship between the non-boundary responses and covariates is well-modelled as linear; see Figure LABEL:fig-supp:xb-data-1 to Figure LABEL:fig-supp:xb-data-4 in the supplementary material document, which show simulated data sets for all parameter settings. The two models start performing similarly again for large $u$, when the response values start concentrating on the boundaries.
@@
Figure 6: Average relative change in $S$ in (18) ($S_{\rm CN}/S_{\rm XBX}-1$) when moving from XBX to CN, based on $100$ samples of size $n=500$ for each combination of $u\in\{2^{-6},2^{-5},\ldots,2\}$, and a subset of $[\mu_{1},\mu_{n}]$ (columns) and $[\phi_{1},\phi_{n}]$ (rows) intervals in Table 4. Positive values of the relative change correspond to $S_{\rm CN}>S_{\rm XBX}$, and, thus, worse predictive performance of the CN model relative to the XBX regression model. The dashed vertical lines are the smallest values of $u$ for which the average probability of boundary observations exceeds $0.95$ for each combination of $[\mu_{1},\mu_{n}]$ and $[\phi_{1},\phi_{n}]$.
The experimental setting in the current section has been used to compare beta regression on rescaled responses according to the proposal in Smithson and Verkuilen (2006) and the normal linear regression model. Figure LABEL:fig-supp:xbx-vs-beta and Figure LABEL:fig-supp:xbx-vs-ols in the Section LABEL:sec:XBX-vs-beta-ols of the supplementary material document show the average relative change in $S$ when moving from XBX regression to each of those models as a function of $u$ for all parameter settings in Table 4. In both cases XBX regression performs best and increasingly better as $u$ increases.
@@ 6 Concluding remarks
We have introduced XBX regression for modelling bounded responses with or without boundary observations. At the core of the XBX regression model is the XB distribution, which is a censored version of a four-parameter beta distribution with the same exceedance $u$ on the left and right of $(0,1)$. We prove that both beta regression with dispersion effects and the heteroscedastic two-limit tobit model are special cases of XB regression, depending on whether $u$ is zero or infinity, respectively. To avoid identifiability issues that may arise in estimating $u$ due to the similarity of the beta and normal distribution for specific parameter settings that result in low skewness and kurtosis, we define XBX regression as a shrunken version of XB regression towards beta regression, by letting $u$ have an exponential distribution with mean $\nu$. The associated marginal likelihood can be conveniently and accurately approximated using a Gauss-Laguerre quadrature rule, resulting in efficient estimation and inference procedures. In Section 5, we produce evidence that XBX regression can substantially outperform heteroscedastic two-limit tobit regression when the common features of skewness and kurtosis of bounded-domain variables are encountered, and perform only slightly worse when skewness and kurtosis are small.
@@
The betareg R package, since version 3.2-0, provides methods for fitting and drawing inference from XBX regression models. Bayesian inference for XBX regression is straightforward through template modelling frameworks such as JAGS (Plummer, 2003) and Stan (Carpenter et al., 2017), due to the alternative definition of XBX regression in Section 3 as the censored version of a mixture of four-parameter beta regressions with the same exponentially-distributed exceedance on the left and right of $(0,1)$. Methods based on Stan are provided by the brms R (Bürkner, 2017) package since version 2.23.